OpenStudy (anonymous):

Solve the inequality 2(x + 5) > 8x – 8. Show your work plz. :)

OpenStudy (anonymous):

You have to use the rules of Algebra.

OpenStudy (anonymous):

\(a(b+c)=ab+ac\) mostly. Then you use properties of inequality.

OpenStudy (anonymous):

ummm. can you show me how to do it

OpenStudy (anonymous):

I can show you an example. \[ \begin{align} 6(x+8)&>10x-10\\ 6x+6 \cdot 8&>10x-10\\ 6x&>10x-10-6 \cdot 8\\ 6x-10x&>-10-6 \cdot 8\\ 6x-10x&>-10-48\\ -4x&>-58\\ x&<\frac{-58}{-4}\\ x&<\frac{29}{2} \end{align} \]

OpenStudy (anonymous):

Apply the same idea to this problem.

OpenStudy (anonymous):

what is the idea...

OpenStudy (anonymous):

sigh...

OpenStudy (anonymous):

Jack, start with distributing the 2 across (x+5), so that side will then look like 2x+ 10

OpenStudy (anonymous):

Then solve for x by getting all the X's to one side, and everything non-X on the other

OpenStudy (anonymous):

So, you'd subtract 8x from the right side, add 8x to the left side, and subtract 10 from both sides, this will isolate your X's to the let side of the equation

OpenStudy (anonymous):

Making it then look like: -6x > -18

OpenStudy (anonymous):

Finally, you have to divide both sides by the -6 to solve for x, because you are dividing by a negative, you have to flip the inequality symbol

OpenStudy (anonymous):

So the answer would be x < 3

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

it helps

OpenStudy (anonymous):

@GREcram, we are not supposed to give full answers. Read the Code of Conduct. I was also going to reply to this, but someone was asking for information. So I will reply now: You do essentially the same as what I did in each step. The only difference is that instead of distributing a \(6\), you distribute a \(2\). Then you subtract the number on the left from the number on the right. Then you subtract the variable on the right from the variable on the left. Then you divide by the coefficient.

OpenStudy (anonymous):

rite your not supposed to give just the answer you are also supposed to explain it. GrEcram did just that

OpenStudy (anonymous):

My bad, I just signed into this site for the first time, like an hour ago.

OpenStudy (anonymous):

@JackJones No, you're not supposed to give the answer. "I will encourage and guide those needing help, and not just give them an answer." See http://openstudy.com/code-of-conduct . @GREcram, it is fine. Simply read the code of conduct. It's honestly not that long.

OpenStudy (anonymous):

she did guide me through the question

OpenStudy (anonymous):

she didnt just type x < 3

OpenStudy (anonymous):

There's more in the URL--which you didn't click, it seems. "Give Help, Not Answers" (this precedes the exact quotation from the URL) You are supposed to strive toward the answer using the help of the members on this site. They aren't supposed to provide you with an answer at any time. They can, however, tell you if you are incorrect and try to send you toward the right path.

OpenStudy (anonymous):

well i didnt understand the problem. that example you gave me was kind of confused

OpenStudy (anonymous):

confusing*

OpenStudy (anonymous):

oh and if your not supposed to only give answers half the ppl on openstudy should be suspended by now

OpenStudy (anonymous):

"oh and if your not supposed to only give answers half the ppl on openstudy should be suspended by now" This isn't a good point. I am aware of this and I agree with you. Most people do not follow the rules of this website. This is a well-documented issue of which the moderators are aware. (Granted, there are special cases where one can give an answer.)

OpenStudy (anonymous):

isn't this a special case

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

i think it is